in two years time,jam will be as old as dino is now.in 26 years dino will be twice as old as jam is now.how old is dino now?
Answer by ankor@dixie-net.com(22740)

in two years time, jam will be as old as dino is now.
therefore
j = d-2
:
in 26 years dino will be twice as old as jam is now.
d + 26 = 2j
replace j with (d-2)
d + 26 = 2(d-2)
d + 26 = 2d - 4
26 + 4 = 2d - d
d = 30 yrs Dino's age now
:
:
Confirm this in the statement:
in 26 years dino will be twice as old as jam is now. (j=28)
30 + 26 = 2(28)
56 = 56